OpenStudy (anonymous):

At time t in seconds, a particle’s distance s(t), in micrometers, from a point is given by s(t)=11sin t + 34. What is the average velocity of the particle from t=π/3 to t=13π/3? (round to 3 decimal places!) would it start off like this? s(13π/3)-s(π/3) ________________ (13π/3 - π/3) ??? :/

jimthompson5910 (jim_thompson5910):

good, keep going

jimthompson5910 (jim_thompson5910):

so you'll have [11*sin(13π/3) + 34] - [11*sin(π/3) + 34] --------------------------------------- 13π/3 - π/3

jimthompson5910 (jim_thompson5910):

evaluate that with a calculator (make sure you are in radian mode)

OpenStudy (anonymous):

i got 39.28180995 so about 39.282 ? is that correct?

jimthompson5910 (jim_thompson5910):

no it's not

OpenStudy (anonymous):

aww okay.. lemme try again.. one sec :)

OpenStudy (anonymous):

i got 0/13π/3 - π/3=-1.047197551 so about -1.047 ? is that right this time? :/

jimthompson5910 (jim_thompson5910):

still not correct

OpenStudy (anonymous):

aww :/ okay.. not sure what i'm doing wrong.. it's in radian mode :/ i'll do it step by step what i get.. one sec:)

OpenStudy (anonymous):

[11*sin(13π/3) + 34] - [11*sin(π/3) + 34] --------------------------------------- 13π/3 - π/3 =(43.52627944) - (43.52627944) --------------------------------- (13.61356817) - (1.047197551) = 0 --------------------- 12.56637062 =0 ? :/

OpenStudy (anonymous):

not sure how i keep getting different values lol :/ hopefully this one's right? :/

jimthompson5910 (jim_thompson5910):

0 is correct

OpenStudy (anonymous):

ohh yay!!! :) 3rd times the charm i guess lol :p thank you!!